# The impact of the redshift-dependent selection effect of halos on the redshift-space power spectrum

### v1

@@ 3.1 The selection effect on real- and redshift-space power spectra for the LOWZ and CMASS mock catalogs
Figure 4: A comparison of the two power spectra, $P_{\rm AM}$ and $P_{M_{\rm th}}$, for the abundance-matching (AM) sample and the single mass-threshold ($M_{\rm th}$) sample both of which reproduce $n(z)$ of the LOWZ sample. We show the ratio for the real-space power spectrum (left panel), and the monopole (middle panel) and quadrupole (right panel) moments of the redshift-space power spectrum. The data points in each panel are the mean of the $50$ realizations and the error bars are the $1\sigma$ error on the mean, estimated by dividing the standard deviation by $\sqrt{50}$. The black dashed vertical line in each panel indicates the half of the Nyquist wavenumber, and the solid line is the ratio computed using the Dark Emulator in Kobayashi et al. [17] (see text for the details).
@@
We now give a theoretical interpretation of the results in Figs. 4 and 5. For this purpose, we use the Dark Emulator developed in Ref. [40] [also see 28], which is a simulation-based emulator that allows a fast, accurate computation of the real- or redshift-space power spectrum of halos for input parameters (cosmological parameters, redshift, and halo masses). For the redshift-space power spectrum, the emulator outputs $P(k,\mu)$, where $\mu$ is a cosine angle between the wavenumber vector and the line-of-sight direction. Hence, we can use the emulator output to compute the monopole and quadrupole moments. The red solid curves in Figs. 4 and 5 show the ratio, $P_{M_{\rm th}}/P_{\rm AM}$, computed using Dark Emulator. We can compute $P_{\rm AM}$ based on Eq. (2.8), substituting the emulator’s output for the local power spectrum,  [^4] For the actual computation of the local power spectrum in Eq. (2.8), we input the cumulative halo number density $\bar{n}(x_{3})$ (Fig. 1) into the emulator, instead of halo mass threshold $M_{\rm th}(x_{3})$ (Fig. 2). while we can compute $P_{M_{\rm th}}$ directly using the emulator prediction for a single mass threshold. We determine this mass threshold so that $P_{M_{\rm th}}$ matches $P_{\rm AM}$ at $k=0.02~{}h\rm Mpc^{-1}$. These emulator predictions fairly well reproduce the simulation result, including the $k$ dependence. Thus, these results provide independent justification for the result that the selection function is small. Please note that the emulator predictions for the monopole and quadrupole moments are about 1% and 5% fractional accuracies for halos of $\sim 10^{13}h^{-1}M_{\odot}$ up to $k\simeq 0.60~{}h{\rm Mpc}^{-1}$ for cosmologies around the fiducial $\Lambda$CDM model, as carefully studied in Ref. [40].
@@ 3.2 The results for halo samples with power-law $\bar{n}(z)$
Figure 6: Similar to Fig. 4, but the results for the power-law samples in the right panel of Fig. 1. The panels in the top, middle and bottom row show the results for $\alpha=0.5,1$ and $2$ of $n(z)\propto z^{-\alpha}$, respectively.

### v2

@@ 3.1 The selection effect on real- and redshift-space power spectra for the LOWZ and CMASS mock catalogs
Figure 4: A comparison of the two power spectra, $P_{\rm AM}$ and $P_{M_{\rm th}}$, for the abundance-matching (AM) sample and the single mass-threshold ($M_{\rm th}$) sample both of which reproduce $n(z)$ of the LOWZ sample. We show the ratio for the real-space power spectrum (left panel), and the monopole (middle panel) and quadrupole (right panel) moments of the redshift-space power spectrum. The data points in each panel are the mean of the 5050 realizations and the error bars are the $1\sigma$ error on the mean, estimated by dividing the standard deviation by $\sqrt{50}$. The black dashed vertical line in each panel indicates the half of the Nyquist wavenumber, and the solid line is the ratio computed using the Dark Emulator in Kobayashi et al. [17] (see text for the details).
@@
[…]
@@
Figure 6: Similarly to Figs. 4 and 5, but here we subtract the residual shot noise from the measured power spectrum for $P_{M_{\rm th}}(k)$ before comparing it with $P_{\rm AM}(k)$. We introduce the residual shot noise parameter as a free parameter for $P_{M_{\rm th}}(k)$ and determine its best-fit value by minimizing the difference between the shot-noise-subtracted $P_{M_{\rm th}}(k)$ and $P_{\rm AM}(k)$ over the range $k=[0.02,0.3]~h{\rm Mpc}^{-1}$. The dashed line in each panel is for the ratio when using $P_{M_{\rm th}}(k)$ with the residual shot noise subtracted, while the solid line is the same as in Figs. 4 and 5.
These upturn behaviors could also be attributed to the effect of residual shot noise [43]. In standard cosmological analyses of the power spectrum [e.g., 16], residual shot noise is commonly included as a nuisance parameter. We determine the residual shot noise by minimizing the the difference between the shot-noise-subtracted $P_{M_{\rm th}}(k)$ and $P_{\rm AM}(k)$ over the range $k=[0.02,0.3]~h{\rm Mpc}^{-1}$. The residual shot noise estimated in this way is: $P_{\rm res}\sim 16$ or $20~(h{\rm Mpc}^{-1})^{3}$ for the LOWZ or CMASS sample, respectively, compared to the naive shot noise $1/\bar{n}_{\rm g}\sim 10^{4}~(h{\rm Mpc}^{-1})^{3}$. We confirm that the upturn behavior in the monopole is reduced if we include the residual shot noise parameter for $P_{\rm M_{\rm th}}$. Note that the quadrupole, or more generally the higher-order moments, are not affected by the residual shot noise.
@@ 3.2 The results for halo samples with power-law $\bar{n}(z)$
Figure 7: Similar to Fig. 4, but the results for the power-law samples in the right panel of Fig. 1. The panels in the top, middle and bottom row show the results for $\alpha=0.5,1$ and 22 of $n(z)\propto z^{-\alpha}$, respectively.